OpenStudy (anonymous):

Convergence test question. Does anybody know how to test if this series converges??

OpenStudy (anonymous):

\[\int\limits_{0}^{0.8} x/2 - x^3/72 +x^5/1440 = 0.39298\]

OpenStudy (zarkon):

'series'?

OpenStudy (anonymous):

I was originally given an integral, then brok it down to these three first non-zero terms

OpenStudy (anonymous):

Then they are asking me to use a test to see if this series converges ...

OpenStudy (anonymous):

So I am guessing I need to put these terms in a formula, or I need to convert them to a series formula ...

OpenStudy (anonymous):

The original integral was \[\int\limits_{0}^{0.8} (1-cost)/t^2 dt \]

OpenStudy (amistre64):

hard to tell what it is your trying to do; but a remainder integration might work if im thinking of it correctly

OpenStudy (anonymous):

I know it converges, but I need to be able to prove why with a convergence test

OpenStudy (anonymous):